OpenStudy (anonymous):

Help please Verify the identity. cos 4x + cos 2x = 2 - 2 sin^2 2x - 2 sin^2 x

OpenStudy (anonymous):

cos 4x = 2cos^2 2x - 1 cos 2x = 2cos^2 x -1 result:2-2sin^ 2x - 2sin^2 x

OpenStudy (anonymous):

so 2cos^2 2x - 1 + 2cos^2 2x - 1= 2-2sin^ 2x - 2sin^2 x?

OpenStudy (asnaseer):

do you know how to express \(\cos(2x)\) in terms of \(\sin(x)\)?

OpenStudy (asnaseer):

using the double-angle formulae

OpenStudy (anonymous):

yes , you should first optimize it wheter you know double angle formulas or not?

OpenStudy (anonymous):

so am I right then? That so 2cos^2 2x - 1 + 2cos^2 2x - 1= 2-2sin^ 2x - 2sin^2 x?

OpenStudy (anonymous):

yes , it was great